OpenStudy (anonymous):

Find the values of a and b that make f continuous everywhere. f(x) = (x2 − 4) / (x − 2) if x < 2 ax2 − bx + 1 if 2 ≤ x < 3 4x − a + b if x ≥ 3

OpenStudy (anonymous):

could you make the question a bit clearer?

OpenStudy (anonymous):

not too sure how..... another one like it tho

OpenStudy (anonymous):

Find the values of a and b that make f continuous everywhere. f(x) = (x2 − 4) / (x − 2) if x < 2 ax2 − bx + 3 if 2 ≤ x < 3 4x − a + b if x ≥ 3 a = b =

myininaya (myininaya):

So you know the left limit needs to equal the right limit for both x=2 and x=3?

myininaya (myininaya):

You have these equations: \[\lim_{x \rightarrow 2^-}f(x)=\lim_{x \rightarrow 2^+}f(x)\] \[\lim_{x \rightarrow 3^-}f(x)=\lim_{x \rightarrow 3^+}f(x)\]

myininaya (myininaya):

Let me know if you need further help. :)

OpenStudy (anonymous):

this is not a limit problem though, it is problem where i need to evaluate both a and b

myininaya (myininaya):

You needs limits here.

myininaya (myininaya):

If the function is continuous at x=2, then you the following things f(2)=limx->2(f(x)) and of course for both sides to exist. In order for limx->2(f(x)) to exist, the left limit needs to equal the right limit.

myininaya (myininaya):

Using what I wrote above, you will get a system of two linear equations involving the unknowns a and b.

OpenStudy (anonymous):

IM SO CONFUSED

myininaya (myininaya):

On the limit part?

OpenStudy (anonymous):

no how to evaluate

OpenStudy (anonymous):

like are my values 2 and 3

myininaya (myininaya):

I looked at where the function was broken up at. I'm trying to make the function exist at those two numbers since those are the only 2 numbers that could possibly make our function discontinuous. If we find the right a and b, we can ensure the function will be continuous at x=2 and x=3. So looking at the limits I asked you to evaluate...we have: \[\lim_{x \rightarrow 2^-}f(x)=\lim_{x \rightarrow 2^+}f(x) \] \[\lim_{x \rightarrow 2^-} \frac{x^2-4}{x-2}=\lim_{x \rightarrow 2^+} (ax^2-bx+1)\] The left hand side can be reduced The right hand side is a polynomial and polynomails are continuous everywhere for any real input. You can just plug in x=2 on the right hand side of the equation. Now the other equation is from \[\lim_{x \rightarrow 3^-}f(x)=\lim_{x \rightarrow 3^+}f(x) \] \[\lim_{x \rightarrow 3^-}(ax^2-bx+1)=\lim_{x \rightarrow 3^+}(4x-a+b) \] Both the left and right expressions are polynomials. Just plug in 3 into both.

OpenStudy (anonymous):

so a(3)^2-b(3)+1 ???

myininaya (myininaya):

So you are looking at the last equation I wrote I assume: a(3)^2-b(3)+1=4(3)-a+b

myininaya (myininaya):

So you have that equation (which you can simplify and combine like terms) Now go back to the first one and evaluate both limits (simplify and combine like terms) You will see we have a system of two linear equations with two unknowns.

myininaya (myininaya):

Did you reduce that fraction yet?

OpenStudy (anonymous):

no i am so confused im sorry

myininaya (myininaya):

\[\frac{x^2-4}{x-2}\] Do you know how to factor x^2-4?

OpenStudy (anonymous):

yes they cancel

myininaya (myininaya):

Yep you reduced the fraction now plug in what x is approaching

myininaya (myininaya):

and you were talking about the (x+2)'s canceling, right?

myininaya (myininaya):

oops the (x-2)'s

OpenStudy (anonymous):

yes so...3? so 3-2=1

myininaya (myininaya):

:( no no... what...where did 3' come from?

myininaya (myininaya):

We are looking at the first equation I wrote.

myininaya (myininaya):

The one where x is approaching 2 not 3.

myininaya (myininaya):

We are done with the approaching 3 equation for now.

myininaya (myininaya):

We will return to it in a bit.

myininaya (myininaya):

So you said (x^2-4)/(x-2)=x+2 right?

myininaya (myininaya):

since the (x-2)'s canceled.

myininaya (myininaya):

Now since x is approaching 2, just plug it in. 2+2=4

myininaya (myininaya):

So that was the left hand side of our equation. Not pluggin 2 on the right hand side of the equation gives you?

OpenStudy (anonymous):

i dont know where to look

myininaya (myininaya):

At the equations I gave you.

myininaya (myininaya):

I gave you two.

myininaya (myininaya):

We already looked at the second. We are looking at the first now.

myininaya (myininaya):

equation 1: \[\lim_{x \rightarrow 2^-} \frac{x^2-4}{x-2}=\lim_{x \rightarrow 2^+} (ax^2-bx+1) \] equation 2: \[\lim_{x \rightarrow 3^-}(ax^2-bx+1)=\lim_{x \rightarrow 3^+}(4x-a+b) \] We already evaluated both sides of equation 2. You were also suppose to simplify and combine like terms for equation 2. Now we are looking back equation 1. We reduced the fraction on the left hand side giving us: \[\lim_{x \rightarrow 2^-} (x+2)=\lim_{x \rightarrow 2^+} (ax^2-bx+1) \] Now since both of these functions are continuous at x=2, plug in x=2.